OpenStudy (anonymous):

How many ordered triples of positive intergers satisfy the equation a+b+c=6

OpenStudy (raden):

eksploration problem case I : for a=1, so b+c=5 or c=5-b c will be an integer positive, if b=1,2,3, and 4 so, the total way = 4 ways case II : for a=2, so b+c=4 or c=4-b c will be an integer positive, if b=1,2, and 3 so, the total way = 3 ways case III : for a=3, so b+c=3 or c=3-b c will be an integer positive, if b=1, and 2 so, the total way = 2 ways case 1V : for a=4, so b+c=2 or c=2-b c will be an integer positive, if b=1 so, the total way = 1 ways the total way from all cases = 4+3+2+1 = 10